Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below.

Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. (Round overhead rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.)

Case A

Case B

Direct materials used $enter a dollar amount (a) $88,000
Direct labor 58,000 150,700
Manufacturing overhead applied 46,400 enter a dollar amount (d)
Total manufacturing costs 147,950 enter a dollar amount (e)
Work in process 1/1/20 enter a dollar amount (b) 18,300
Total cost of work in process 204,800 enter a dollar amount (f)
Work in process 12/31/20 enter a dollar amount (c) 13,400
Cost of goods manufactured 193,500 enter a dollar amount (g)
Answer #1
Amount $
Case A Case B
Direct materials used     43,550     88,000
Direct labor     58,000 150,700
Manufacturing overhead applied     46,400 120,560
Total manufacturing costs 147,950 359,260
Work in process 1/1/20     56,850     18,300
Total cost of work in process 204,800 377,560
Work in process 12/31/20     11,300     13,400
Cost of goods manufactured 193,500 364,160
Working:
a.)     43,550 =147950-58000-46400
b.)     56,850 =204800-147950
c.)     11,300 =204800-193500
Manufacturing overhead rate 80.00% =46400/58000
d.) 120,560 =150700*80.00%
e.) 359,260 =88000+150700+120560
f.) 377,560 =359260+18300
g.) 364,160 =377560-13400